# Ultra-fast Digital DPC Yielding High Spatio-Temporal Resolution for Low-Dose Phase Characterisation

### v1

@@ Acknowledgements
Figure S1: Peak Signal to Noise Ratio (PSNR) in finite dose iCOM images. Cut-out of the top left corner of the reconstructed images for detector geometries annular four-segmented, 4x4 pixel array, and 16x16 pixel array are shown beneath.
Figure S2: PSNR in |COM| images for finite doses $6241$, $1500$, and $624.1$ e/Å2, plotted with the same y-axis in order to clearly show how the PSNR goes down with decreasing dose. The reference image is the infinite dose |COM| image as calculated from the full 4D-STEM
@@
Figure S5: Unfiltered and high-pass filtered images and their difference. The high-pass filtered image was computed by subtracting a low-pass filtered version of the unfiltered image (Gaussian blur with $\sigma=10$).

### v2

@@ Acknowledgements
S1 4D-STEM detectors
Summarised in Table S1 are numbers representing the dynamic ranges of the commercially available 4D-STEM detectors displayed in Figure 1 of the paper. While it is inadvisable to compare the dynamic range of detectors directly using either of the given numbers, they are important to take into account because certain experiments may require data with a higher dynamic range. The bit depth gives information about how many bits is used to read out from the detector the information contained in each frame. Note that bit depth does not necessarily tell the full story of the dynamic range of a detector since bits may be used to store other information than the intensity values directly. Additionally, Table S1 gives the highest beam current a detector can withstand without saturating in electrons per pixel per second. This is a physical property of the detector, however it can not be used to compare detectors directly because it does not account for their differing number of pixels.
Table S1: Detector types and dynamic range, given in bit depth and maximum incident beam current before saturation in e-/pixel/s.
Detector Type Bit depth Maximum count rate (e-/pixel/s)
EMPAD PAD 30 $1.2\cdot 10^{7}$
EMPAD2 PAD 32 $10^{9}$
Arina PAD 8-321118-32 image bit depth, 12 read-out bit depth. $10^{8}$
Ela PAD 8-16222Speed given in the plot for the Dectris Ela is for the 8 bit read-out. $6.2\cdot 10^{6}$
UltraScan CCD 16 -
Orius CCD 14 -
OneView MAPS 12 $2.25\cdot 10^{5}$
Rio Scintillator CMOS 12 $6\cdot 10^{5}$
K2 MAPS 32 $10^{*}$
K3 MAPS 32333The K3 reads out at 32 bit in linear accumulation mode, and 8 bit in counting mode. $40^{*}$
Metro MAPS 32 $80^{*}$
ClearView MAPS 12 $3\cdot 10^{6}$
Medipix 1 PAD 15 -
MerlinEM PAD 1-24444For the Medipix/MerlinEM detectors that support a large range of bit depths the speed given in the plot is for 12 bit, but the acquisition speed may be increased by using a lower bit depth. $10^{6}$
Medipix 4 PAD 1-24 -
pnCCD CCD 16 $4\cdot 10^{5}$
Celeritas MAPS - $8.7\cdot 10^{5}$
DE-16 MAPS - $10^{5}$
Recommended maximum dose rate to ensure detector linearity.
S2 COM with varying detector geometries
Figure S1: Peak Signal to Noise Ratio (PSNR) in simulated finite dose iCOM images. Cut-out of the top left corner of the reconstructed images for detector geometries annular four-segmented, 4x4 pixel array, and 16x16 pixel array are shown beneath.
Figure S2: PSNR in simulated |COM| images for finite doses $6241$, $1500$, and $624.1$ e/Å2, plotted with the same y-axis in order to clearly show how the PSNR goes down with decreasing dose. The reference image is the infinite dose |COM| image as calculated from the full 4D-STEM
@@
Figure S5: Unfiltered and high-pass filtered experimental images and their difference. The high-pass filtered image was computed by subtracting a low-pass filtered version of the unfiltered image (Gaussian blur with $\sigma=10$).
Figure S6: Screenshot of the software controlling the Pulse digitiser unit. The display shows the raw data stream (in orange), the gradient (in blue), the threshold (horisontal gray line), and the digitised events (in green) for Channel 1. The sidebar shows the settings for this channel.